OpenStudy (anonymous):

Solve the following system of equations -3x + 5y = 68 x + 4y =51

OpenStudy (anonymous):

i think....pick any eq for finding another eq. ex: x+4y=51...so bring 4y to 51 so the eq will be like this.....x=51-4y.....then sub it to -3x+5y=68.....that u can find y....and sub the y to the eq x+4y=51...then u get x....

OpenStudy (anonymous):

-3x+5y=68 x+4y=51 solving by cross multiplying x/(5*51-68*4)=y/(68*1-51*-3)=(-1)/(-3*4-5*1) x/(255-272)=y/(68+153)=(-1)/(-12-5) x/(-17)=y/221=(-1)/█(-17@.) x/(-17)=(-1)/█(-17@.), Or x=(-1)/(-17)*-17=-1 y/221=(-1)/(-17) Y=(-1)/(-17)*221 Or y=13

OpenStudy (anonymous):

thanks